OpenStudy (anonymous):

Help please i need to solve this with derivatives A ball is thrown vertically upward from a height of 6 ft with an initial velocity of 60 ft per second. How high will the ball go?

OpenStudy (anonymous):

174.75

OpenStudy (zehanz):

General formula for anobject with a stating speed \(v_0\) formula that is subjected to an accelleration a is: \(v_t=v_0+at\). Use g instead of a (it is a negative number, because it tries to slow the object down) and solve \(v_t=0\) for t.

OpenStudy (anonymous):

those formulas are also with derivatives?

OpenStudy (zehanz):

Then put the result in the formula for the distance covered during accellerated movement: \(S_t=v_0t=\dfrac{1}{2}at^2\), again using g instead of a (and with a negative sign)

OpenStudy (anonymous):

I have to solve it with calculus and derivatives though, thats physics right?

OpenStudy (zehanz):

Typo: \(S_t=v_0t+\frac{1}{2}at^2\) As you can see, the derivative of this formula is the first formula.

OpenStudy (zehanz):

YOu don't have to use calcuus, if you just use the formulas. Then it is algebra only.

OpenStudy (anonymous):

the thing is that i have to solve it with calculus since its for my calculus class

OpenStudy (zehanz):

OK, you can do that as follows: First calculate how long it takes to get to zero speed: \(v_t=v_0+at=60-32.185t=0\). Call the solution \(t_1\). Then the distance covered is the integral of the speed: \[S=\int\limits_{0}^{t_1}v(t)dt=\int\limits_{0}^{t_1}(60-32.185t)dt\]

OpenStudy (zehanz):

32.185 being the accelleration due to gravity, in ft/s²

OpenStudy (anonymous):

ouuuu okokkk

OpenStudy (zehanz):

YW! BTW, if you do the same with the formula \(v(t)=v(0)t+at\), you get: \[S(t)=\int\limits_{0}^{t}v(0)+a \tau d \tau=v(0)t+\frac{1}{2}at^2\], that is the formula I suggested you at the start.